If you have a 108 square feet then enclose it by sqaure what is the length and width?

Answer by stanbon(75887)

If you have a 108 square feet then enclose it by
square what is the length and width?
----
Length = Width = sqrt(108) = sqrt(36*3) = 6sqrt(3)ft.
====================
Cheers,
Stan H.
====================